OpenStudy (anonymous):

-2=4(s+8)-3s. what is s?

OpenStudy (anonymous):

-2=4s+32-3s, -34=s

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

first step distribute 4 to s+8. so -2=4s+32-3s. subtract 32, so -2-32=4s-3s. combine like terms so -34=s